Write the equation of the line passing through the points ( negative 2, 7) and (3, negative 5)
Answer by algebrahouse.com(1659)

"Write the equation of the line passing through the points (-2, 7) and (3,-5)"

One way to get the equation of a line,
is to find the slope and y-intercept.
Then, substitute those into the slope-intercept form: y = mx + b


(-2,7) , (3,-5)
y2 - y1
-------- {formula to find slope}
x2 - x1

-5 - 7
-------- {substituted into slope formula}
3 - (-2)

-12
--- {simplified}
5

slope is -12/5

Now, pick one of the points, and substitute it with the slope, into slope-intercept form.
This will give you the y-intercept.

(3,-5), m = -12/5
y = mx + b {slope-intercept form}
-5 = -12/5(3) + b {substituted into slope-intercept form}
-5 = -36/5 + b {multiplied -12/5 by 3}
b = 11/5 {added 36/5 to both sides}
y-intercept is 11/5

Substitute slope and y-intercept into slope-intercept form y = mx + b
m is the slope
b is the y-intercept

y = -12/5x + 11/5
www.algebrahouse.com